# Chatbot Companionship: A Mixed-Methods Study of Companion Chatbot Usage Patterns and Their Relationship to Loneliness in Active Users

### v1

@@ 4.3 Multiple Regression Model
Figure 11: The final model from our exploratory multiple regression, depicting the relationships between chatbot session length, loneliness, and related factors. While session length did not directly predict loneliness, neuroticism and problematic use of chatbots were positively associated with loneliness, while social network size was negatively associated. Agreeableness also showed a negative association with loneliness. One significant interaction effect was found: session length interacted with social attraction to a close person.
[…]
The analysis also revealed a slight but significant interaction effect (see Figure 12) of session length and social attraction to a close person ($\beta=0.054$, $p=0.049$), suggesting that social attraction to a close person is a moderator between the relationship between session length and loneliness. Social attraction to the chatbot approached significance ($\beta=0.0800$, $p=0.061$), while social attraction to a close person, age, and sex did not significantly predict loneliness in this model.
@@
The mediation analysis revealed nuanced relationships between chatbot session length, loneliness, and the proposed mediators (Figure 13). The total effect of chatbot session length on loneliness was positive but not statistically significant ($\beta=0.0797$, 95% CI $[-0.0034,0.1612]$). The direct effect, controlling for mediators, was also positive but not significant ($\beta=0.0606$, 95% CI $[-0.0196,0.1386]$). Examining the indirect effects, we found that problematic internet use (GPIUS2) significantly mediated the relationship between chatbot session length and loneliness ($\beta=0.0183$, 95% CI $[0.0027,0.0403]$), suggesting that increased chatbot usage is associated with higher problematic internet use, which in turn is related to increased loneliness. However, social network size (LSNS) did not show a significant mediating effect ($\beta=0.0007$, 95% CI $[-0.0304,0.0340]$). The total indirect effect, while positive, was not statistically significant ($\beta=0.0191$, 95% CI $[-0.0149,0.0563]$).
@@
Figure 13: Results of bootstrapped parallel multiple mediation analysis (5,000 resamples) examining the mediating effects of social network size (LSNS) and problematic chatbot use (GPIUS2) on the relationship between chatbot session length and loneliness.
@@
Figure 14: Model summary of an ordinary least squares regression for the effect of chatbot session length on loneliness (ULS-8), using standardized variables. Model Formula: Loneliness (ULS-8) $\sim$ Session Length Score + Social Attraction towards Chatbot + Socialization (LSNS) + Social Attraction towards Close Person + Neuroticism + Sex + Age + Session Length Score $\times$ Social Attraction towards Close Person
4.4 Clusters of users of companion chatbots
@@
Our K-means cluster analysis revealed seven distinct user profiles among companion chatbot users, each exhibiting unique behaviors, motivations, and psychological characteristics. The resulting clusters are visualized in the heatmap and dendrogram presented in Figure 15. Detailed heatmaps including all numerical scales can be seen in Figure 16. A version with the original values of the scales can be found in Figure 17 in Appendix B.
@@
Validation of the cluster solution using Kruskal-Wallis tests revealed significant differences among clusters in terms of session length, loneliness (ULS-8), problematic use of chatbots (GPIUS-2), and frequency of use (all $p<0.001$). The average silhouette score ($-0.0413$) suggested some overlap between clusters, which is not uncommon in complex psychological data. A MANOVA test confirmed significant multivariate differences among the clusters when considering all variables simultaneously (Wilks’ $\lambda=0.2420,F(24,1375.71)=28.77,p<0.001$). All four MANOVA test statistics (Wilks’ lambda, Pillai’s trace, Hotelling-Lawley trace, and Roy’s greatest root) were significant ($p<0.001$), providing robust evidence for cluster differences across multiple dimensions. Box’s M-test indicated significant heterogeneity of covariance matrices ($\chi^{2}=181.96$, $p<0.001$), which does not invalidate the clustering but suggests the need for cautious interpretation. This heterogeneity may affect the reliability of post-hoc comparisons and could indicate that the relationship between variables differs across clusters. While these results provide strong support for the distinctiveness of the identified clusters, they should still be considered exploratory, providing a foundation for more targeted investigations in future research.
Despite these statistical considerations, the 6-cluster solution provided meaningful and actionable insights for our study context, balancing statistical rigor with practical interpretability. This approach aligns with the exploratory nature of cluster analysis in psychological research, where the goal is often to identify patterns for further investigation [Hair et al.(2010)]. Future research should aim to replicate these cluster profiles in independent samples and explore the stability of these clusters over time.
@@
4.4.1 Cluster 0: Disengaged Light Users
4.4.2 Cluster 1: Well-Adjusted Moderate Users
4.4.3 Cluster 2: AI-Wary Light Users
4.4.4 Cluster 3: Lonely Moderate Users
4.4.5 Cluster 4: Fulfilled Dependent Users
4.4.6 Cluster 5: Lonely Light Users
4.4.7 Cluster 6: Socially Challenged Frequent Users
@@ 5 Discussion
Our study provides a detailed examination of the relationship between companion chatbot usage and loneliness, offering some insights into the complex interplay of psychological, social, and technological factors that shape this relationship. Through a combination of regression analysis and cluster identification, we developed a nuanced understanding of how different user groups engage with and are affected by companion chatbots.
5.1 A Model for Understanding the Relationship Between Companion Chatbot Usage and Loneliness
Our regression analysis revealed a complex relationship between companion chatbot usage and loneliness. While session length did not have a significant direct relationship with loneliness, the presence of a significant interaction term with social attraction towards a close person suggests a more nuanced, potentially nonlinear relationship.
We differentiated between mediators and confounders based on theoretical considerations. Potential mediators, such as problematic use of chatbots, social attraction towards the chatbot, and socialization, were identified as variables that could reasonably be affected by chatbot usage. Confounders, like personality traits, were considered as factors likely to influence both usage and loneliness but less likely to be affected by usage in the short term.
Our final model identified problematic use as a significant mediator in the relationship between chatbot usage and loneliness. This finding aligns with intuitive expectations; increased usage provides more opportunities for problematic behaviors to develop. The positive association between usage and problematic use suggests that as individuals spend more time with chatbots, they may be at higher risk of developing patterns of use that could contribute to feelings of loneliness.
While not significant in terms of mediation effects, the negative association between usage and socialization (specifically, social network size) warrants attention. This relationship resonates with concerns about technology potentially replacing human interactions. On one side, this relationship might suggest that as individuals spend more time with chatbots, they may allocate less time to maintaining and expanding their human social networks, potentially contributing to increased loneliness. However, the reverse causality is also possible, with people turning to chatbots to fill a social gap when they cannot or do not interact with other people in their lives.
The relationship between usage and social attraction towards the chatbot, while not a significant mediator, is logically consistent. More time spent interacting with the chatbot may allow for deeper attachments to form. However, the direction of causality between this attraction and loneliness remains unclear. Social attraction to the chatbot could lead to loneliness due to unfulfilled relational needs, or loneliness could potentially enhance social attraction to the chatbot if users seek validation and attention from this artificial source.
Among the confounding variables, neuroticism showed a strong positive association with loneliness, while agreeableness demonstrated a strong negative association. These findings align with previous research on personality traits and loneliness [Buecker et al.(2020)], reinforcing the importance of controlling for these factors in studies of technology use and social well-being.
@@ 5.1 A Model for Understanding the Relationship Between Companion Chatbot Usage and Loneliness
While our model provides valuable insights into the associations between different factors, it is important to acknowledge its limitations. The exploratory nature of our model selection process warrants cautious interpretations and further investigation into the relationships it implies. The assumption of linear relationships may not fully capture the complexity of the relationship between chatbot usage and loneliness. Additionally, while we can theorize about the directions of relationships between variables, our cross-sectional design limits causal inferences. Future research employing experimental designs or longitudinal studies could provide more definitive evidence of causal relationships and explore potential non-linear effects. We plan to follow up with our participants to see changes in psychological, social, and usage characteristics at different time points.
@@ 5.2 Understanding Clusters of Chatbot Users
Our cluster analysis revealed seven distinct user profiles, each with unique patterns of chatbot usage, loneliness levels, and associated psychological and social characteristics.
The contrast between the Socially Fulfilled Dependent Users (Cluster 4) and the Lonely Moderate Users (Cluster 3) is striking. Both groups exhibit high problematic use, yet they differ dramatically in their loneliness levels and overall well-being. This disparity underscores the importance of considering the broader context of an individual’s social life when evaluating the impact of chatbot use.
Socially Fulfilled Dependent Users demonstrate that it is possible to engage extensively with chatbots while maintaining high levels of well-being. Their high empathy and attraction towards chatbots, when considered in conjunction with their more neutral or even negative attitudes towards other people, average socialization, and high self-esteem, suggest a high level of emotional engagement that is socially fulfilling. The positive relational effects reported by this group, such as increased ease in talking to people due to interacting with chatbots, indicate that chatbot use may even enhance their social skills and confidence.
In contrast, Lonely Moderate Users show high levels of problematic use and loneliness, coupled with low socialization, perceived social support, and self-esteem. Their high preference for chatbots over humans suggests that excessive chatbot use may be contributing to social withdrawal. Again, we note that the direction of causality is unclear – pre-existing loneliness and social difficulties may have led to increased chatbot use, rather than the reverse.
The Well-Adjusted Moderate Users (Cluster 1) present an interesting middle ground. Like the Socially Fulfilled Dependent Users, they exhibit low loneliness and high perceived social support. However, their more moderate chatbot usage and lower emotional connection to AI suggest that they prioritize human connections while still finding value in chatbot interactions, perhaps using them more as tools than as social companions.
@@
The AI-Wary Light Users (Cluster 2) offer yet another perspective on chatbot engagement. These users show low emotional attachment to chatbots and low opinions of AI in general. This group may be using chatbots purely for practical purposes, without seeking or developing emotional connections. Many traits typically associated with loneliness – neuroticism, agreeableness, socialization, and self-esteem) are fairly average, but their loneliness is the second lowest of all clusters. Considering this in comparison to other clusters suggests that, while companion chatbots might be an effective solution to loneliness for Fulfilled Dependent Users, not using companion chatbots or remaining emotionally distant from them may be better for others in terms of their loneliness.
The existence of these diverse user profiles highlights the complex and multifaceted nature of human-AI interactions. It suggests that the relationship between companion chatbot usage, loneliness, and social well-being is not uniform but highly dependent on individual characteristics, usage patterns, and broader social contexts.
@@ 5.3 Implications and Future Directions
Our findings have several implications for the design, deployment, and societal impact of companion chatbots:
5.3.1 Potential Benefits and Risks

### v2

@@ 4.2.2 Multiple Regression Model
A parallel multiple mediation analysis (Figure 12) revealed that problematic internet use (GPIUS2) significantly mediated the relationship between session length and loneliness ($\beta=0.0183$, 95% CI $[0.0027,0.0403]$), while social network size (LSNS) did not show significant mediation ($\beta=0.0007$, 95% CI $[-0.0304,0.0340]$). The total indirect effect was not statistically significant ($\beta=0.0191$, 95% CI $[-0.0149,0.0563]$).
@@
Figure 12: Results of bootstrapped parallel multiple mediation analysis (5,000 resamples) examining the mediating effects of social network size (LSNS) and problematic chatbot use (GPIUS2) on the relationship between chatbot session length and loneliness, including lower and upper confidence intervals (CI_Lower and CI_Upper).
@@
Figure 13: Model summary of an ordinary least squares regression for the effect of chatbot session length on loneliness (ULS-8), using standardized variables. Model Formula: Loneliness (ULS-8) $\sim$ Session Length Score + Social Attraction towards Chatbot + Problematic Use (GPIUS2) + Socialization (LSNS) + Social Attraction towards Close Person + Neuroticism + Agreeableness + Sex + Age + Session Length Score $\times$ Social Attraction towards Close Person
4.3 RQ3: Can we identify and characterize distinct user profiles based on usage patterns, loneliness levels, problematic use, and other psychological and social factors?
4.3.1 Overall characteristics
Participants ranged in age from 18 to 73, primarily 30-35. Most (93.6%) were from the United States, white/Caucasian (66.3%), and employed full-time (61.6%), with a broad distribution of household income. For relationship status, most were single (42.1%) or married (32.4%). 59.7% of participants identified as men and 36.6% identified as women. Additional demographic characteristics are in Figure 14.
Most participants utilize Snapchat MyAI (31.6%), Character.ai (30.4%), or Replika (22.6%) and tend to spend 15 and 30 minutes in each session with a companion chatbot (42.6%), with 35.4% reported sessions shorter than 15 minutes. Very few users spend extended periods chatting, with only 1.28% having sessions exceeding 4 hours. Most participants have been using their companion chatbots for a few months (43.9%) to around a year (35.2%), with only 7.7% of participants having used their chatbots for multiple years and 13.2% for around a month or less. More details on usage can be seen in Figure 4 and Figure 5.
Figure 14: Demographic table for study participants, including overall percentage for each category and the difference in percentage from the overall percentage for each cluster, with darker colors indicating a greater difference. Blue: cluster has a lower percentage of the category compared to overall. Red: cluster has a higher percentage of the category compared to overall.
4.3.2 Clusters of users of companion chatbots
@@
Our K-means cluster analysis revealed seven distinct user profiles among companion chatbot users, each exhibiting unique behaviors, motivations, and psychological characteristics. The resulting clusters are visualized in the heatmap and dendrogram presented in Figure 15. The demographics of each cluster compared to overall demographics can be seen in Figure 14, with percentages found in Figure 20 of Appendix B.
@@
Validation of the cluster solution using Kruskal-Wallis tests revealed significant differences among clusters in terms of session length, loneliness (ULS-8), problematic use of chatbots (GPIUS2), and frequency of use (all $p<0.001$). The average silhouette score ($-0.0413$) suggested some overlap between clusters, which is not uncommon in complex psychological data. A MANOVA test confirmed significant multivariate differences among the clusters when considering all variables simultaneously (Wilks’ $\lambda=0.2420,F(24,1375.71)=28.77,p<0.001$). All four MANOVA test statistics (Wilks’ lambda, Pillai’s trace, Hotelling-Lawley trace, and Roy’s greatest root) were significant ($p<0.001$), providing robust evidence for cluster differences across multiple dimensions. Box’s M-test indicated significant heterogeneity of covariance matrices ($\chi^{2}=181.96$, $p<0.001$), which does not invalidate the clustering but suggests the need for cautious interpretation. This heterogeneity may affect the reliability of post-hoc comparisons and could indicate that the relationship between variables differs across clusters. While these results provide strong support for the distinctiveness of the identified clusters, they should still be considered exploratory, providing a foundation for more targeted investigations in future research.
@@
The seven identified clusters represent distinct patterns of chatbot engagement and psychological characteristics:
  • 0: Disengaged Light Users (7.67%): Minimal chatbot engagement with average loneliness levels. Show neutral attitudes toward AI technology and average social support levels. Primary usage is brief (<15 min) casual conversations, with main motivations being curiosity to start use and practical utility for continued use. Average age is 32; for gender, 65% are men and 35% are women.
  • 1: Well-Adjusted Moderate Users (23.02%): Average usage with low loneliness. Demonstrate high extraversion, low neuroticism, and strong social networks. Show positive attitudes toward both AI and human interactions. Main motivations of curiosity to start use and emotional support to continue use. Average age is 36; 67% men, 33% women.
  • 2: AI-Wary Light Users (11.88%): Average to low usage with low loneliness. Show high skepticism toward AI technology while maintaining strong human connections. Users were interested in chatbots for practical utility and recreation, with continued use motivated by practical purposes, such as seeking knowledge and advice. Average age is 34; 65% men, 35% women.
  • 3: Lonely Moderate Users (13.61%): Average usage with high loneliness. Show low socialization but positive attitudes toward chatbots. Notable for seeking emotional support and companionship through chatbot interactions, with some participants integrating chatbots into their lifestyle. Average age is 35; 51% men, 45% women, 4% other.
  • 4: Fulfilled Dependent Users (18.32%): High usage with below-average loneliness. Show strong emotional engagement with chatbots while maintaining adequate social connections. Report positive impacts on real-world social interactions. Average age is 35; 61% men, 37% women, 1% other.
  • 5: Lonely Light Users (14.60%): Average usage with high loneliness. Demonstrate high neuroticism and low self-esteem. Users seek out chatbots for emotional support and companionship, often discussing personal issues and sometimes engaging in intimate relationships with chatbots. Average age is 34; 47% men, 52% women.
  • 6: Socially Challenged Frequent Users (10.89%): Average frequency but short sessions, with high loneliness. Show low extraversion and social support, with notably low trust in both familiar and unfamiliar people. Both initial and continued use is motivated by a desire for companionship, with a few rare cases engaging in intimate relationships with their chatbot. Casual exchange was a less common theme of conversation. Average age is 36; 61% men, 39% women.
Detailed heatmaps including all numerical scales can be seen in Figure 16. A version with the original values of the scales can be found in Figure 19 along with detailed cluster descriptions in Appendix B. The patterns of themes for each cluster’s motivations for use and topics of discussion can be seen in Figure 6 and Figure 7.
Post-hoc analysis using Dunn’s test with Bonferroni correction revealed distinct patterns among clusters across our key variables (Figure 17). For session length, Fulfilled Dependent Users (Cluster 4) showed significantly longer sessions than all other clusters (p < .001) except Well-Adjusted Moderate Users. Disengaged Light Users (Cluster 0) had significantly shorter sessions than most other clusters (p < .001).
Usage frequency showed less variation between clusters, with only Disengaged Light Users showing significantly lower frequency than all other clusters (p < .001), and Fulfilled Dependent Users showing higher frequency than AI-Wary Light Users and Lonely Light Users (p < .01).
For loneliness (ULS-8), three clusters - Lonely Moderate Users, Lonely Light Users, and Socially Challenged Frequent Users (Clusters 3, 5, and 6) - showed significantly higher levels than Well-Adjusted Moderate Users and AI-Wary Light Users (p < .001). Fulfilled Dependent Users maintained significantly lower loneliness levels than these high-loneliness clusters (p < .001).
Problematic use (GPIUS2) was highest in Lonely Moderate Users and Fulfilled Dependent Users (Clusters 3 and 4), who showed significantly higher levels than all other clusters (p < .001) except each other. AI-Wary Light Users (Cluster 2) demonstrated the lowest levels of problematic use (p < .001).
Figure 17: Significant pairwise differences between clusters on key variables based on Dunn’s test with Bonferroni correction. Cell background color indicates significance of differences (darker green = presence of greater significance in differences), with text color indicating direction (red = greater than, blue = less than). For example, >2*** indicates significantly greater values than Cluster 2 at p < .001. Significance levels: *: p < .05, **: p < .01, ***: p < .001.
@@ 5 Discussion
Our study provides a detailed examination of the relationship between companion chatbot usage and loneliness, offering some insights into the complex interplay of psychological, social, and technological factors that shape this relationship.
5.1 Understanding the Relationship Between Companion Chatbot Usage and Loneliness
To answer RQ2, “What is the relationship between companion chatbot usage and loneliness, and what factors mediate or moderate this relationship?” we look to our regression analysis. While session length showed no significant direct relationship with loneliness, our regression analysis revealed a more nuanced picture through mediating and moderating factors. Problematic use emerged as a significant mediator, suggesting that increased usage may create more opportunities for problematic behaviors to develop. This aligns with existing research connecting problematic internet use as both a cause and effect of loneliness [Kim et al.(2009)], indicating a similar pattern may exist for companion chatbot use.
The negative association between usage and social network size warrants attention, though causality remains unclear. Users might reduce human interactions as they spend more time with chatbots, or they might turn to chatbots to fill an existing social void. The positive relationship between usage and social attraction towards chatbots is logically consistent – more interaction time allows deeper attachments to form. However, whether this attraction leads to loneliness through unfulfilled relational needs or loneliness enhances attraction to chatbots is uncertain.
Among confounding variables, neuroticism showed a strong positive association with loneliness, while agreeableness demonstrated a strong negative association, aligning with previous research on personality traits and loneliness [Buecker et al.(2020)], reinforcing the importance of controlling for these factors in studies of technology use and social well-being.
@@ 5.2 Understanding Clusters of Chatbot Users
We shall first discuss the overall characteristics of our participants, which we explored for RQ1, “What motivates the use of companion chatbots, and what do people use them for?” Our analysis revealed that despite companion chatbots’ marketing as social tools, most participants primarily used them for technological exploration and entertainment rather than friendship or intimate relationships. This usage pattern may parallel other solitary activities like gaming or reading, which can affect socialization and lead to overindulgence. However, our cluster analysis revealed distinct user profiles demonstrating more complex relationships with these technologies. This brings us to RQ3, “Can we identify and characterize distinct user profiles based on usage patterns, loneliness levels, problematic use, and other psychological and social factors?”
The most striking contrast emerges between Socially Fulfilled Dependent Users (Cluster 4) and Lonely Moderate Users (Cluster 3). Both exhibit high problematic use, yet their outcomes differ dramatically. Socially Fulfilled Dependent Users maintain high well-being despite (or perhaps because of) their extensive chatbot engagement. Their high empathy and attraction towards chatbots, combined with positive reported effects on real-world social interactions, suggests that intensive chatbot use can coexist with healthy social functioning. These findings parallel work by Snodgrass et al. showing that more intensively involved gamers, those exhibiting what appears to be “addictive” play, experience the greatest psychosocial benefits from their gaming – including reduced loneliness [Snodgrass et al.(2018)]. However, unlike gaming where more intense involvement tends to predict better social outcomes, we found that Lonely Moderate Users with similar usage patterns experienced markedly different results. For chatbots, the nature of engagement may matter more than intensity alone. This challenges concerns about the negative effects of AI replacing human relationships [Pataranutaporn et al.(2021), Turkle(2011)] and merits further investigation into how these users maintain this balance.
Conversely, Lonely Moderate Users demonstrate how similar usage patterns can lead to markedly different outcomes. Their high loneliness, low socialization, and preference for chatbots over humans raise concerns about whether chatbot use exacerbates social withdrawal or merely reflects existing social difficulties. Their pattern of emotional disclosure to chatbots suggests these tools may serve as their primary outlet for emotional expression. We suggest further investigation of whether chatbots have a positive or negative effect on these individuals.
Socially Challenged Frequent Users (Cluster 6) and Lonely Light Users (Cluster 5) are another interesting pair to consider. Socially Challenged Frequent Users are most similar to Lonely Moderate Users (3), though they have average problematic use and considerably below-average measures of trust and empathy, even for those they feel close to. These individuals likely need intervention beyond what a companion chatbot can currently provide, yet companion chatbots may be the only entities they are willing to engage with. Both Socially Challenged Frequent Users and Lonely Light Users seek out chatbots for companionship, and it is largely only in these two groups that participants discussed having or seeking a romantic or sexual relationship with their chatbot. The participants in these two clusters most strongly reflect the Media Equation, in that their responses to chatbots may be similar to their social responses to people [Reeves and Nass(1998)]. Of concern is whether or not Lonely Light Users might transition to Socially Challenged Frequent Users over time, losing trust and empathy with people as they engage more with chatbots. Considering that loneliness is a risk factor for problematic use [Kim et al.(2009)], understanding how these users’ characteristics can change over time, and how chatbot systems can detect these at-risk individuals and intervene, may be valuable for future research.
Well-Adjusted Moderate Users (Cluster 1), the largest group, demonstrate a potentially optimal engagement pattern. With moderate usage and lower emotional connection to AI, they maintain strong human connections while leveraging chatbots primarily as tools for emotional processing rather than social companions. This suggests that chatbots might effectively reduce loneliness without necessarily functioning as relationship substitutes. Many disclose their emotional issues to chatbots, valuing the safe space for expression they provide, with a few considering the chatbot as part of their lifestyle. Chatbots might effectively reduce loneliness without necessarily functioning as relationship substitutes. Further investigation into causality and mechanisms of psychosocial changes may prove fruitful for understanding how companion chatbots can be designed to reduce loneliness, besides the typical strategy of pushing for greater emotional intimacy with the user. For consideration, recent work has shown that social support acts as a significant mediator between psychological problems and problematic gaming [Malak et al.(2023)], which may help explain how Cluster 4, with poor perceived social support, exhibits high loneliness – as opposed to Cluster 1, with high perceived social support and low loneliness – when both have moderate chatbot usage.
@@
The AI-Wary Light Users (Cluster 2) provide an important counterpoint, showing that emotional distance from chatbots can coexist with low loneliness. These users show low emotional attachment to chatbots and low opinions of AI in general. This group uses chatbots mostly for practical purposes and recreation, perhaps without seeking or developing emotional connections. Many traits associated with loneliness – neuroticism, agreeableness, socialization, and self-esteem) are fairly average, but their loneliness is the second lowest of all clusters. Considering this in comparison to other clusters suggests that emotional investment in AI companions may not be necessary or even optimal for all users.
These diverse profiles highlight that the impact of companion chatbot use on loneliness and social well-being varies significantly based on individual characteristics and usage patterns. This heterogeneity suggests the need for personalized approaches in both chatbot design and potential intervention strategies.
@@ 5.3 Implications and Future Directions
Our findings extend beyond simple social displacement theories [Kraut et al.(1998)]suggesting a framework where companion chatbots serve different social functions based on user characteristics and usage patterns. This aligns with and extends the Parasocial Contact Hypothesis [Schiappa et al.(2005)], which suggests that media characters can provide meaningful social experiences. Chatbot interactions may similarly provide parasocial benefits, but with important distinctions based on user traits and usage patterns.
The mediating role of problematic use in the relationship between usage and loneliness contributes to theories of technology addiction and compensatory internet use [Kardefelt-Winther(2014)]. Excessive chatbot use may reflect attempts to compensate for social deficits, similar to problematic social media use, but with unique characteristics due to the interactive nature of companion chatbots. This extends existing frameworks for understanding problematic technology use to include AI companionship specifically.
5.3.1 Theoretical Contributions